OpenStudy (smarty):

Describe the nature of the roots of the equation 2x^2 + 3x = –3.

OpenStudy (anonymous):

\[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] a=2 b=3 c=3 subtitution and solve

OpenStudy (anonymous):

imaginary numbers

OpenStudy (smarty):

one real root two real roots two imaginary roots cannot be determined

OpenStudy (smarty):

So C? two imaginary roots?

OpenStudy (smarty):

@Nancy_Lam

OpenStudy (anonymous):

yeah

OpenStudy (smarty):

Thanks~

OpenStudy (smarty):

what about: 7x^2 = 4x + 1.

OpenStudy (smarty):

@sritama

OpenStudy (anonymous):

two imaginary roots

OpenStudy (smarty):

How can you tell?

OpenStudy (smarty):

For ex: x^2 – 4x + 4 = 0. how would you tell for that one?

OpenStudy (anonymous):

for the 2nd equation it is b

OpenStudy (anonymous):

7x^2 = 4x + 1 , i mean

OpenStudy (smarty):

haha alright what about : x^2 – 4x + 4 = 0

OpenStudy (anonymous):

two real roots,and they are same

OpenStudy (anonymous):

x^2 – 4x + 4 = 0 (x-2)^2

OpenStudy (anonymous):

all you have to get from the value of \[\sqrt{b^2-4ac}\]

OpenStudy (smarty):

Ohok. thanks!

OpenStudy (smarty):

So the root of hte discriminant.

OpenStudy (anonymous):

if the value is negatives, roots are imaginary if positive . roots are real if 0,roots are real and same

OpenStudy (anonymous):

yeah